3. The tank has 700 liters oil and is being drainad ar a constart rate of 25 liters per minute.




a) How do you write a linear function V for the number of liten in the tank after t minutes




Tassuming that the drainage started art-012




bi Compute how many litera are in the tank after 75 minutes and 55 seconds.


1
Expert's answer
2021-11-22T13:27:55-0500

a)


"V(t)=700-25t"

b)


"75\ min\ 55 sec=(75+\dfrac{55}{60})min=\dfrac{911}{12}\ min"

"V(t)=700-25(\dfrac{911}{12})<0"

We cannot consider the time 75min 55 sec.

How long will it take for the tank to be drained empty?


"V=700-25t=0"

"t=\dfrac{700}{25}"

"t=28\ min"